Solve by factoring.

1. 3a squared - 36a + 81 = 0
2. 5x squared - 45 = 0
3. 2a squared - a - 21 = 0
4. 3n squared - 11n + 10 = 0
5. 2x squared - 7x -9 = 0

Found 2 solutions by checkley71, edjones:
Answer by checkley71(8403)

3A^2-36A+81=0
(3A-9)(A-9)=0 ANSWER.
5X^2-45=0
5(X^2-9)=0
5(X+3)(X-3)=0 ANSWER.
2A^2-A-21=0
2A-7)(A+3)=0 ANSWER.
I'LL LEAVE THE REST FOR YOU TO PRACTICE WITH.

Answer by edjones(7311)

1.
3a^2-36a+81=0
3*81=243 multiply the 1st and last coefficients.
we need 2 neg. numbers that when multiplied=243 and when added=-36
-9, -27 are factors of 243 that add to -36. Finding these factors is the only hard part.
(3a^2-9a)+(-27a+81)=0 Factor in parts. Notice (3a^2-27a)+(-9a+81)=0 gives the same answer.
3a(a-3)-27(a-3)=0
(3a-27)(a-3)=0
3(a-9)(a-3)=0
a=9, a=3
.
4.
3n^2-11n+10=0
3*10=30
We need 2 negative numbers that when multiplied=30 and when added=-11
-5 and -6 fill the bill.
(3n^2-6n)+(-5n+10)=0
3n(n-2)-5(n-2)=0
(3n-5)(n-2)=0
3n-5=0
3n=5
n=5/3, n=2
.
2.
5x^2-45=0
5(x^2-9)=0
5(x+3)(x-3)=0
x=-3, x=3
.
I did the 3 toughest.
The other 2 are similar to the 1st 2 I did.
I'll leave them up to you.
.
Ed